OpenStudy (anonymous):

Help and you get a medal and fan!! In the last baseball game, there were two 3-run home runs and four hits that each scored two runs. Write an algebraic expression to model the total score in each situation. The evaluate then expression to find the total score. I think it would be something like 2*3 + 4x +2?? I really have no idea, the question makes no sense

OpenStudy (anonymous):

well if x = # of 3-run home runs and y = # of hits that scored 2 runs, then we have 2x + 4y since we know there were 2 home runs and 4 hits, we can say x = 2 and y = 4 and plug them in 2x+4y 2*2 + 4*4 ... plug in x = 2 and y = 4 4 + 16 20 so the total score is 20

OpenStudy (anonymous):

ok that seems to make sense...one question...shouldnt the 2x be 3x? since it was a 3-run home run? or no?

OpenStudy (anonymous):

@GabbyCakes

OpenStudy (anonymous):

@B_O_R_E_D Oh, oh okay. I think its like this: the expression should be 3x + 2y then we plug in x = 2 and y = 4 to get 3x + 2y 3*2 + 2*4 6 + 8 14 so the total score is 14

OpenStudy (anonymous):

Ok I see what you did....That makes sense...thank you! :)

OpenStudy (anonymous):

Your welcome! Sorry for the quick mess up. I was in a moment of insanity cx

OpenStudy (anonymous):

lol no biggie

OpenStudy (anonymous):

2x +3y then try * 3*2 =6 what you do to one you do to all so 2*4 =8 8+6=14